# A Comprehensive Survey of Datasets for Clinical Mental Health AI Systems

### v1

@@ Abstract
Mental health disorders are rising worldwide. However, the availability of trained clinicians has not scaled proportionally, leaving many people without adequate or timely support. To bridge this gap, recent studies have shown the promise of Artificial Intelligence (AI) to assist mental health diagnosis, monitoring, and intervention. However, the development of efficient, reliable, and ethical AI to assist clinicians is heavily dependent on high-quality clinical training datasets. Despite growing interest in data curation for training clinical AI assistants, existing datasets largely remain scattered, under-documented, and often inaccessible, hindering the reproducibility, comparability, and generalizability of AI models developed for clinical mental health care. In this paper, we present the first comprehensive survey of clinical mental health datasets relevant to the training and development of AI-powered clinical assistants. We categorize these datasets by mental disorders (e.g., depression, schizophrenia), data modalities (e.g., text, speech, physiological signals), task types (e.g., diagnosis prediction, symptom severity estimation, intervention generation), accessibility (public, restricted or private), and sociocultural context (e.g., language and cultural background). Along with these, we also investigate synthetic clinical mental health datasets. Our survey identifies critical gaps such as a lack of longitudinal data, limited cultural and linguistic representation, inconsistent collection and annotation standards, and a lack of modalities in synthetic data. We conclude by outlining key challenges in curating and standardizing future datasets and provide actionable recommendations to facilitate the development of more robust, generalizable, and equitable mental health AI systems.
@@ Introduction
The collection and dissemination of clinical mental health data is hampered by stringent privacy regulations, such as the General Data Protection Regulation (GDPR) [1] in the European Union and the Health Insurance Portability and Accountability Act (HIPAA) [2] in the United States. These frameworks place strict limitations on the use and sharing of sensitive health information, complicating the creation of large, representative datasets that meet ethical and legal standards. Early data-driven approaches often turned to social media platforms such as Reddit [3, 4], Twitter [5, 6], and YouTube [7, 8] to compile mental health datasets [9]. While these sources offer scale and accessibility, they generally lack clinical validation and may not accurately reflect diagnosed populations, limiting their clinical applicability.
@@
In recent years, researchers have begun constructing clinically grounded datasets through collaborations with mental health institutions, often under strict privacy-preserving protocols and with informed consent. These datasets, derived from electronic health records, clinician notes, therapy transcripts, and physiological or neuroimaging data, are often shared only upon request or remain private, following extensive pseudonymization procedures. Despite increased interest in such resources, there is no comprehensive review focused specifically on clinical mental health datasets for AI development. Previous reviews have addressed broader applications of AI in mental health [10, 11], or specific methodological areas such as machine learning [12], deep learning [13, 14], natural language processing (NLP)[15, 16], and large language models (LLMs)[17, 18].
@@
Although some recent surveys (see Table 1) have incorporated data-centric perspectives, they often treat datasets as secondary to model performance or application domains. Reviews have examined resources spanning electronic health records, imaging, social media, wearables, and digital platforms [19, 20, 21, 22, 23], or focused on natural language processing [24, 25, 26], yet few provide a systematic analysis of dataset characteristics. Key considerations, such as accessibility (public, restricted, private), cultural and linguistic representation, and the emerging role of synthetic data, remain largely unaddressed. In contrast, our survey provides a dedicated and comprehensive examination of clinical mental health datasets for AI development. We categorize datasets by disorder type, access level, task formulation, data modality, and sociocultural context, then identify gaps that limit model generalizability and clinical relevance. We further highlight current efforts in synthetic data generation to address privacy and data scarcity. By shifting the focus from algorithms to data foundations, our work offers actionable guidance for curating and evaluating datasets that support robust, ethical, and equitable mental health AI systems (see Figure 2).
@@ 40 FEBD
Many psychiatric symptoms manifest across multiple behavioral channels, making multimodal analysis essential for accurate diagnosis. For depression, relevant cues span speech semantics [163], prosody [164, 165, 166, 167], and facial expressions [168]. The AVEC-2017 [169] and AVEC-2019 [170] challenges demonstrated that models incorporating text, audio, and visual information outperform unimodal baselines in depression severity prediction. Similarly, the AVEC-2018 challenge [171] showed improved performance in bipolar disorder diagnosis through multimodal fusion.
The most common combination is text and audio, seen in 22 clinical datasets targeting conditions like depression, anxiety, schizophrenia, bipolar disorder, and PTSD [56, 57, 65, 66, 67, 68, 69, 70]. These datasets typically comprise interview transcripts and audio recordings from clinical or semi-structured interactions. Some, like MMPsy [58] and Jeong et al.[71], collect responses to structured prompts. Following a similar approach, Ex-Ray [76] records participants speaking and writing a passage about a particular topic with specific target words. Other datasets embed diagnostic tasks, including Theory of Mind assessments[72, 172, 173], social role-play [73], and narrative responses to TAT images [74, 174]. Other tasks include phoneme task, testing phonemic and semantic fluency, picture description and prompted narrating [75]. Hayati et al. [77] collect and transcribe semi-structured interviews for depression detection in different Malay dialects.
@@
Federated learning (FL) can address the scale and diversity limits of mental health datasets but remains vulnerable to gradient leakage attacks [187, 188, 189, 190]. Combining FL with local differential privacy (LDP) [190, 191] adds protection by locally injecting noise, yet often degrades performance on small clinical datasets [192]. Advancing secure, effective deployment requires novel LDP-FL methods that better balance privacy and utility, validated in realistic, resource-constrained settings.

### v2

# A Comprehensive Review of Datasets for Clinical Mental Health AI Systems

@@ Abstract
Mental health disorders are rising worldwide. However, the availability of trained clinicians has not scaled proportionally, leaving many people without adequate or timely support. To bridge this gap, recent studies have shown the promise of Artificial Intelligence (AI) to assist mental health diagnosis, monitoring, and intervention. However, the development of efficient, reliable, and ethical AI to assist clinicians is heavily dependent on high-quality clinical training datasets. Despite growing interest in data curation for training clinical AI assistants, existing datasets largely remain scattered, under-documented, and often inaccessible, hindering the reproducibility, comparability, and generalizability of AI models developed for clinical mental health care. In this paper, we present the first comprehensive review of clinical mental health datasets relevant to the training and development of AI-powered clinical assistants. We categorize these datasets by mental disorders (e.g., depression, schizophrenia), data modalities (e.g., text, speech, physiological signals), task types (e.g., diagnosis prediction, symptom severity estimation, intervention generation), accessibility (public, restricted or private), and sociocultural context (e.g., language and cultural background). Along with these, we also investigate synthetic clinical mental health datasets. We identify critical gaps such as a lack of longitudinal data, limited cultural and linguistic representation, inconsistent collection and annotation standards, and a lack of modalities in synthetic data. We conclude by outlining key challenges in curating and standardizing future datasets and provide actionable recommendations to facilitate the development of more robust, generalizable, and equitable mental health AI systems.
@@ Introduction
The collection and dissemination of clinical mental health data is hampered by stringent privacy regulations, such as the General Data Protection Regulation (GDPR) [1] in the European Union and the Health Insurance Portability and Accountability Act (HIPAA) [2] in the United States. These frameworks place strict limitations on the use and sharing of sensitive health information, complicating the creation of large, representative datasets that meet ethical and legal standards. Early data-driven approaches often turned to social media platforms such as Reddit [3], Twitter [4], and YouTube [5] to compile mental health datasets [6]. While these sources offer scale and accessibility, they generally lack clinical validation and may not accurately reflect diagnosed populations, limiting their clinical applicability.
@@
In recent years, researchers have begun constructing clinically grounded datasets through collaborations with mental health institutions, often under strict privacy-preserving protocols and with informed consent. These datasets, derived from electronic health records, clinician notes, therapy transcripts, and physiological or neuroimaging data, are often shared only upon request or remain private, following extensive pseudonymization procedures. Despite increased interest in such resources, there is no comprehensive review focused specifically on clinical mental health datasets for AI development. Previous reviews have addressed broader applications of AI in mental health [7], or specific methodological areas such as machine learning [8], deep learning [9], natural language processing (NLP)[10], and large language models (LLMs)[11].
@@
Although some recent review articles (see Table 1) have incorporated data-centric perspectives, they often treat datasets as secondary to model performance or application domains. Reviews have examined resources spanning electronic health records, imaging, social media, wearables, and digital platforms [12, 13, 14, 15, 16], or focused on natural language processing [17, 18, 19], yet few provide a systematic analysis of dataset characteristics. Key considerations, such as accessibility (public, restricted, private), cultural and linguistic representation, and the emerging role of synthetic data, remain largely unaddressed. In contrast, our review provides a dedicated and comprehensive examination of clinical mental health datasets for AI development. We categorize datasets by disorder type, access level, task formulation, data modality, and sociocultural context, then identify gaps that limit model generalizability and clinical relevance. We further highlight current efforts in synthetic data generation to address privacy and data scarcity. By shifting the focus from algorithms to data foundations, our work offers actionable guidance for curating and evaluating datasets that support robust, ethical, and equitable mental health AI systems (see Figure 2).
To identify relevant resources for this survey, we conducted a systematic search across Google Scholar, DBLP, and the ACL Anthology, using a set of targeted keywords covering both general and disorder-specific domains. These included terms such as “Mental Health AI Datasets,” “Multimodal Mental Health AI Datasets,” “Mental Health AI Datasets for Depression,” “Multimodal Mental Health AI Datasets for Depression,” “Mental Health AI Datasets for Anxiety,” “Multimodal Mental Health AI Datasets for Anxiety,” “Mental Health AI Datasets for Bipolar Disorder,” “Multimodal Mental Health AI Datasets for Bipolar Disorder,” “Mental Health AI Datasets for PTSD,” “Multimodal Mental Health AI Datasets for PTSD,” “Mental Health AI Datasets for Schizophrenia,” “Multimodal Mental Health AI Datasets for Schizophrenia,” “MRI Mental Health AI Datasets,” and “EEG Mental Health AI Datasets”. This search initially yielded $560$ records, which, after deduplication and screening for relevance, were narrowed down to $89$ datasets that met our selection criteria of containing real-world clinical mental health data applicable to AI research.
@@ 40 FEBD
Many psychiatric symptoms manifest across multiple behavioral channels, making multimodal analysis essential for accurate diagnosis. For depression, relevant cues span speech semantics [156], prosody [157], and facial expressions [158]. Evidence from prior works further supports the efficacy of multimodal approaches. The AVEC-2017 [159] and AVEC-2019 [160] challenges demonstrated that models incorporating text, audio, and visual information outperform unimodal baselines in depression severity prediction. Similarly, the AVEC-2018 challenge [161] showed improved performance in bipolar disorder diagnosis through multimodal fusion.
The most common combination is text and audio, seen in 22 clinical datasets targeting conditions like depression, anxiety, schizophrenia, bipolar disorder, and PTSD [49, 50, 58, 59, 60, 61, 62, 63]. These datasets typically comprise interview transcripts and audio recordings from clinical or semi-structured interactions. Some, like MMPsy [51] and Jeong et al.[64], collect responses to structured prompts. Following a similar approach, Ex-Ray [69] records participants speaking and writing a passage about a particular topic with specific target words. Other datasets embed diagnostic tasks, including Theory of Mind assessments[65], social role-play [66], and narrative responses to TAT images [67]. Other tasks include phoneme task, testing phonemic and semantic fluency, picture description and prompted narrating [68]. Hayati et al. [70] collect and transcribe semi-structured interviews for depression detection in different Malay dialects.
@@
Federated learning (FL) can address the scale and diversity limits of mental health datasets but remains vulnerable to gradient leakage attacks [174]. Combining FL with local differential privacy (LDP) [174] adds protection by locally injecting noise, yet often degrades performance on small clinical datasets [175]. Advancing secure, effective deployment requires novel LDP-FL methods that better balance privacy and utility, validated in realistic, resource-constrained settings.